OpenStudy (anonymous):

Solve: 2(x + 1) = −8

OpenStudy (anonymous):

2x +2 = -8 -2 -2 x= -6

OpenStudy (anonymous):

Thanks but these are the options 5 −3 −5 −4.5

OpenStudy (anonymous):

hmm very weird.

OpenStudy (anonymous):

Its five

ganeshie8 (ganeshie8):

2(x + 1) = −8 you need to isolate x, so perform the opposite operation divide both sides by 2

ganeshie8 (ganeshie8):

x+1 = -8/2 x+1 = -4

ganeshie8 (ganeshie8):

perform the opposite operation again, subtract both sides -1

ganeshie8 (ganeshie8):

x+1-1 = -4-1 x = -5

OpenStudy (anonymous):

Never mind its -5

ganeshie8 (ganeshie8):

cool :)

OpenStudy (anonymous):

Yeah but thanks =)

ganeshie8 (ganeshie8):

yw :)